C3H8+5O2->3CO2+4H2O
Based on the reactions, how many molecules of H2O are produced when 2 moles of O2 are reacted with an excess of propane.
1
Expert's answer
2020-06-22T09:17:03-0400

From the equation we have that 1 molecule of C3H8 reacts with 5 molecules of O2 to produce 3 molecules of CO2 and 4 molecules of H2O,

Ratio of O2 used to H2O produced is 5:4

Now, if we use 2 molecules of O2, then molecules of H2O produced is given by,

(2×4)/5=8/5=1.6 moles of H2O are produced.